# Machine Learning Fleet Efficiency: Analyzing and Optimizing Large-Scale Google TPU Systems with ML Productivity Goodput

### v1

@@ 5. Improving Fleetwide Efficiency
Looking at the aggregated MPG of the fleet does not necessarily help ML practitioners identify what kinds of improvements will make the largest impact on the fleet; this is where the decomposability of the metric comes into play, as shown in Table 2. By breaking MPG into its three components; Program Goodput, Runtime Goodput, and Scheduling Goodput, we can diagnose fleetwide issues and identify the types of optimizations that would most improve fleet efficiency. Furthermore, we can segment the fleet using the characteristics described in Section 3 in order to identify issues with specific workload types and propose model-level optimizations.
@@
5.1. Program Goodput Optimizations
Example: Quantifying the impact of an XLA optimization on the TPU fleet. It is rare for any single optimization to have a significant impact on overall fleet-wide PG. But we can track the impact of these optimizations by looking at the change in PG for a fixed set of benchmarked workloads or segment of the production fleet over time. For example, looking at a benchmark of the top 150 most costly workloads in the fleet, Figure 12 pinpoints the effect of a code change that was submitted to the XLA compiler - in this case, an algebraic simplification in the compiler graph. The dramatic increase in PG for the benchmark of 150 workloads suggests that the positive impact of this optimization can be generalized to the ML fleet as a whole.
It is also helpful to look at PG fluctuations across hardware segments of the fleet. Figure 13 illustrates a notional example where looking at segmented PG can uncover insights that would otherwise be hidden by looking at aggregate metrics. In this case, the segmented data suggests that when a new ML accelerator chip is introduced to the fleet, the workloads running on that chip may initially have a low PG, since the model / compiler code has not been fully tailored for that chip yet. As user adoption increases and accelerator-specific software optimizations are rolled out to the fleet, PG gets closer to theoretical peak efficiency. In other words, hardware accelerator maturity tends to yield greater PG over time. As the chip nears the end of its lifecycle (represented by decreasing allocation in the fleet, and illustrated in Figure 13 by the “Chip decommissioned” label), the PG decreases due to lower chip usage and natural workload/compiler drift. This highlights the importance of co-design across all layers of the ML fleet to make sure that both the software (compiler) and hardware (chips) are optimized for the latest workloads.
@@ 5.2. Runtime Goodput Optimizations
5.3. Scheduling Goodput Optimizations

### v2

@@ 5. Improving Fleetwide Efficiency
Looking at the aggregated MPG of the fleet does not necessarily help ML practitioners identify what kinds of improvements will make the largest impact on the fleet; this is where the decomposability of the metric comes into play, as shown in Table 2. By breaking MPG into its three components; Scheduling Goodput, Runtime Goodput, and Program Goodput, we can diagnose fleetwide issues and identify the types of optimizations that would most improve fleet efficiency. Furthermore, we can segment the fleet using the characteristics described in Section 3 in order to identify issues with specific workload types and propose model-level optimizations.
@@
5.1. Scheduling Goodput Optimizations
@@ 5.2. Runtime Goodput Optimizations
5.3. Program Goodput Optimizations
Example: Quantifying the impact of an XLA optimization on the TPU fleet. It is rare for any single optimization to have a significant impact on overall fleet-wide PG. But we can track the impact of these optimizations by looking at the change in PG for a fixed set of benchmarked workloads or segment of the production fleet over time. For example, looking at a benchmark of the top 150 most costly workloads in the fleet, Figure 11 pinpoints the effect of a code change that was submitted to the XLA compiler - in this case, an algebraic simplification in the compiler graph. The increase in PG for the benchmark of 150 workloads suggests that the positive impact of this optimization can be generalized to the fleet.
It is also helpful to look at PG fluctuations across hardware segments of the fleet. Figure 15 shows a notional example where looking at segmented PG can uncover insights that would otherwise be hidden by looking at aggregate metrics. In this case, the segmented data suggests that when a new ML accelerator chip is introduced to the fleet, the workloads running on that chip may initially have a low PG, since the model / compiler code has not been fully tailored for that chip yet. As user adoption increases and accelerator-specific software optimizations are rolled out to the fleet, PG gets closer to theoretical peak efficiency. In other words, accelerator maturity tends to yield greater PG over time. As the chip nears the end of its lifecycle (represented by decreasing allocation in the fleet, and shown in Figure 15 by the “Chip decommissioned” label), the PG decreases due to lower chip usage and natural workload/compiler drift. This shows the importance of co-design across all layers of the ML fleet to make sure that both the software (compiler) and hardware (chips) are optimized for the latest workloads.
6. Discussion
To put things into perspective for MPG, we considered more traditional approaches to measuring ML fleet efficiency such as Capacity, Occupancy, and Duty Cycle. Figure 16 illustrates how we may have traditionally tended to think about performance metrics (Li et al., 2023; Mars et al., 2011; Kanev et al., 2015). However, our experience reveals that these traditional metrics can sometimes fall short in providing a holistic view, given the unique challenges we have discussed in Section 2 and Section 3. While we cannot provide specific measured details for proprietary reasons, we can share the key insights and lessons learned from our evaluation of these metrics in a production environment.
High Capacity equates to high resource availability. While capacity can tell us how many individual accelerators may be available in the fleet at a given time, it does not take into account the topological shape of those accelerators. For example, an ML training workload requesting thousands of chips in a certain physical mesh shape may never be scheduled if the only available accelerators are fragmented across different clusters or data centers. Other factors, such as the geographical location of data storage cells and accelerators, are not included in the capacity metric, even though they significantly affect job scheduling. Therefore, high capacity by itself as a metric does not necessarily effectively translate to high availability for workloads, and we should instead opt for scheduling efficiency as a more robust metric.
High Occupancy guarantees productivity. Occupancy is defined as the fraction of accelerators allocated to jobs and is often measured by the scheduler (e.g. Borg (Verma et al., 2015)). Occupancy is traditionally seen as a key efficiency indicator, but it can be misleading as it masks inefficiencies in the system stack. For example, an accelerator might be successfully allocated but stuck in I/O wait or running poorly optimized code, thus resulting in a high occupancy but very little actual progress being made towards the workload task. This is important for long-running tasks such as ML model training, where frequent pre-emptions may hinder checkpoint progress but still result in a nominally high occupancy. The traditional occupancy metric therefore does not distinguish between productive and unproductive use of allocated resources.
Duty Cycle accurately represents useful work. Duty Cycle measures whether an accelerator is in use, not how much of its compute capacity is used. When looking at an ML workload running on a TPU, duty cycle does not provide any signal on how much the matrix-multiply units (MXUs) are utilized (Jouppi et al., 2023). It is agnostic of the program-level efficiency and does not take into account the effectiveness of the operations being performed. An accelerator could have a high duty cycle while executing unnecessary or redundant computations. So, we require a more sophisticated metric.
Figure 16. Traditional utilization-based metrics. We replace these using goodput as a measure of fleet efficiency rather than utilization.
Overarching Misconception: Utilization == Productivity. The thread among these metrics is the assumption that keeping accelerators busy equates to productive work. However, this overlooks critical factors. (1) Quality of Computations: None of these metrics assess whether the operations being performed are actually contributing to the desired output. (2) Workload Efficiency: They do not consider whether the workloads are optimally designed for the hardware. (3) System-level Bottlenecks: Focusing solely on accelerator usage ignores potential bottlenecks in data loading, memory access, or inter-accelerator communication. (4) Forward Progress: The traditional metrics provide no insight into how much useful work is being accomplished towards completing an actual ML task, thus motivating the need for a goodput-based metric, MPG.